How much 20% salt solution is added to 20 liters of 60% salt solution to get a solution that is 36% salt
Answer by RAY100(1637)

Let x= amt of added solution
.
.2x + .6*20 = .36(20+x) = 7.2 + .36x
.
4.8 = .16x
.
30 = x,,,,,in liters
.
.
check,,.2(30) + .6(20) = .36(50) = 6+12 = 18,,,,,,ok
.